# Quantum metrology with a continuous-variable system

### v1

@@ Acknowledgments
This work was supported by the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+. This work was funded by MCIN/AEI/10.13039/501100011033 and the European Union ‘NextGenerationEU’ PRTR fund [RYC2021-031094-I], by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA Project call—QUANTUM SPAIN Project, by the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). This work was supported through the project CEX2023-001292-S funded by MCIU/AEI.
@@ Appendix A Basic identities
$1\displaystyle 2\langle\xi,0|\hat{n}|0,\xi\rangle+$ 1 $=2\sinh^{2}r+1=\cosh(2r)$ (159)
$\sinh(2r)$ $=\sqrt{\cosh^{2}(2r)-1}=2\sqrt{n(n+1)}\;.$ (160)
@@ Appendix B First moment of quadrature operators
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$ $\dfrac{\sqrt{2}\text{Im}[\alpha]\sin[\gamma]}{e^{2\left|\alpha\right|^{2}}+\cos[\gamma]}$ $-\dfrac{\sqrt{2}\text{Re}[\alpha]\sin[\gamma]}{e^{2\left|\alpha\right|^{2}}+\cos[\gamma]}$
Compass $\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$ 00 00
@@ Appendix D First and second moments of the number operator
Gaussian state $(\mu^{2}+|\nu|^{2})n_{T}+|\alpha|^{2}+|\nu|^{2}$ $\left(2|\nu|^{4}+8\mu^{2}|\nu|^{2}+2\mu^{4}\right)n_{T}^{2}$ $\quad+\left(4|\alpha|^{2}|\nu|^{2}+3|\nu|^{4}-2\alpha^{2}\mu\nu^{*}-2(\alpha^{*})^{2}\mu\nu+4|\alpha|^{2}\mu^{2}+8\mu^{2}|\nu|^{2}+\mu^{4}\right)n_{T}$ $\quad+\left(|\alpha|^{4}+3|\alpha|^{2}|\nu|^{2}+|\nu|^{4}-\alpha^{2}\mu\nu^{*}-(\alpha^{*})^{2}\mu\nu+|\alpha|^{2}\mu^{2}+2\mu^{2}|\nu|^{2}\right)$
@@
Fock superposition $(|m\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$ $\dfrac{m+n}{2}$ $\dfrac{m^{2}+n^{2}}{2}$
@@
Compass $\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$ $\left|\alpha\right|^{2}\frac{1-e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\sin(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}$ $\left|\alpha\right|^{2}\frac{1-e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\sin(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}+\left|\alpha\right|^{4}\frac{1+e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}$
@@ Appendix E Wigner functions
$W_{mn}(x,p)=\begin{cases}\frac{1}{\pi}(-1)^{m}\left(\frac{n!}{m!}\right)^{1/2}e^{-(x^{2}+p^{2})}\left(-\sqrt{2}(x-ip)\right)^{m-n}L_{n}^{m-n}\left(2(x^{2}+p^{2})\right)\quad&\text{if }m\geq n\\ \frac{1}{\pi}(-1)^{n}\left(\frac{m!}{n!}\right)^{1/2}e^{-(x^{2}+p^{2})}\left(-\sqrt{2}(x+ip)\right)^{n-m}L_{m}^{n-m}\left(2(x^{2}+p^{2})\right)&\text{if }m<n\end{cases}.$ (187)
@@
$\begin{split}W_{0n}(x,p)&=\frac{e^{-(x^{2}+p^{2})}}{2\pi}\left(\frac{2^{n/2}}{\sqrt{n!}}\left((x+ip)^{n}+(x-ip)^{n}\right)+(-1)^{n}L_{n}\left(2\left(p^{2}+x^{2}\right)\right)+1\right)\;.\end{split}$ (188)
@@
$W_{\beta}(x,p)=\dfrac{\left(e^{-((x-\sqrt{2}\beta)^{2}+(p-\sqrt{2}\beta_{i})^{2})}+e^{-((x+\sqrt{2}\beta_{r})^{2}+(p+\sqrt{2}\beta_{i})^{2})}+2e^{-(x^{2}+p^{2})}\cos(2\sqrt{2}(x\beta_{r}-p\beta_{i})-\gamma)\right)}{2\pi\left(1+e^{-2\left|\beta\right|^{2}}\cos\gamma\right)}\;.$ (189)
@@
Appendix F Gaussian quantum Fisher information for symplectic transformations
@@ F.1 General expression
Here we derive the QFI for a single-mode Gaussian state under a quadratic Hamiltonian. The general form of the QFI for Gaussian states is given in Eq. (29). Using $\frac{\partial}{\partial\theta}\mathcal{S}(\theta)=\Omega H\mathcal{S}(\theta)$ in Eq. (II.3), we obtain
$\frac{\partial}{\partial\theta}\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta)}$ $=\Omega H\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta)},$
@@
This leads to
$\left(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right)^{2}$ $=\left(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]-H\Omega\right)^{2}$
$=\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]-H\Omega\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]$
$\quad-\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]H\Omega+H\Omega H\Omega,$ (191)
with trace
@@
$\frac{1}{2}\,\mathrm{Tr}\{\left(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right)^{2}\}$ $=\,\mathrm{Tr}\{\Omega H\Omega H\}-\,\mathrm{Tr}\{\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]H\Omega\}$
$=\,\mathrm{Tr}\{\Omega H\Omega H\}-\,\mathrm{Tr}\{\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\Omega\},$ (192)
where in the last step we used $\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]=(\mathcal{S}(\theta)^{T})^{-1}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\mathcal{S}(\theta)^{-1}$.
The Gaussian QFI for quadratic unitary evolutions described by the Hamiltonian matrix $H$ thus is independent of $\theta$ and reads
$F_{Q}[\hat{\rho},\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}]=\frac{\,\mathrm{Tr}\{\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\}+\,\mathrm{Tr}\{\Omega H\Omega H\}}{1+(4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}])^{-1}}+\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}.$ (193)
Notice further that for a $2\times 2$ matrix ${A}$ the following identities hold:
@@
$A^{-1}$ $=\frac{1}{\det{A}}\left[(\,\mathrm{Tr}{A}){I}-{A}\right]$
$(\,\mathrm{Tr}{A}){I}-{A}$ $=\Omega A^{T}\Omega^{T},$ (194)
and therefore

### v2

@@ Acknowledgments
This work is supported by the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+, by the project CNS2024-154818 with funding by MICIU/AEI /10.13039/501100011033, by the project RYC2021-031094-I, with funding from MCIN/AEI/10.13039/501100011033 and the European Union ‘NextGenerationEU’ PRTR fund, by the project CIPROM/2022/66 with funding by the Generalitat Valenciana, and by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA Project call—QUANTUM SPAIN Project, by the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). This work is supported through the project CEX2023-001292-S funded by MCIU/AEI.
@@ Appendix A Basic identities
$2\langle\xi,0|\hat{n}|0,\xi\rangle+1$ $=2\sinh^{2}r+1=\cosh(2r)$ (165)
$\sinh(2r)$ $=\sqrt{\cosh^{2}(2r)-1}=2\sqrt{\overline{n}(\overline{n}+1)}\;.$ (166)
@@ Appendix B First moment of quadrature operators
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$ $\dfrac{\sqrt{2}\text{Im}[\alpha]\sin[\gamma]}{e^{2\left|\alpha\right|^{2}}+\cos[\gamma]}$ $\dfrac{\sqrt{2}\text{Re}[\alpha]\sin[\gamma]}{e^{2\left|\alpha\right|^{2}}+\cos[\gamma]}$
Compass $\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$ 0 0
@@ Appendix D First and second moments of the number operator
Gaussian state $(\mu^{2}+|\nu|^{2})n_{T}+|\alpha|^{2}+|\nu|^{2}$
$\left(2|\nu|^{4}+8\mu^{2}|\nu|^{2}+2\mu^{4}\right)n_{T}^{2}$
$\quad+\left(4|\alpha|^{2}|\nu|^{2}+3|\nu|^{4}-2\alpha^{2}\mu\nu^{*}-2(\alpha^{*})^{2}\mu\nu+4|\alpha|^{2}\mu^{2}+8\mu^{2}|\nu|^{2}+\mu^{4}\right)n_{T}$
$\quad+\left(|\alpha|^{4}+3|\alpha|^{2}|\nu|^{2}+|\nu|^{4}-\alpha^{2}\mu\nu^{*}-(\alpha^{*})^{2}\mu\nu+|\alpha|^{2}\mu^{2}+2\mu^{2}|\nu|^{2}\right)$
@@
Fock superposition
$(|m\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$
$\dfrac{m+n}{2}$ $\dfrac{m^{2}+n^{2}}{2}$
@@
Compass
$\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$
$\left|\alpha\right|^{2}\frac{1-e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\sin(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}$ $\left|\alpha\right|^{2}\frac{1-e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\sin(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}+\left|\alpha\right|^{4}\frac{1+e^{-2\left|\alpha\right|^{2}}-2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}{1+e^{-2\left|\alpha\right|^{2}}+2e^{-\left|\alpha\right|^{2}}\cos(\left|\alpha\right|^{2})}$
@@ Appendix E Wigner functions
$W_{mn}(x,p)=\begin{cases}\frac{1}{\pi}(-1)^{m}\left(\frac{n!}{m!}\right)^{1/2}e^{-(x^{2}+p^{2})}\left(-\sqrt{2}(x-ip)\right)^{m-n}L_{n}^{m-n}\left[2(x^{2}+p^{2})\right]\quad&\text{if }m\geq n\\ \frac{1}{\pi}(-1)^{n}\left(\frac{m!}{n!}\right)^{1/2}e^{-(x^{2}+p^{2})}\left(-\sqrt{2}(x+ip)\right)^{n-m}L_{m}^{n-m}\left[2(x^{2}+p^{2})\right]&\text{if }m<n\end{cases}.$ (193)
@@
$\begin{split}W_{0n}(x,p)&=\frac{e^{-(x^{2}+p^{2})}}{2\pi}\left(\frac{2^{n/2}}{\sqrt{n!}}\left((x+ip)^{n}+(x-ip)^{n}\right)+(-1)^{n}L_{n}\left[2\left(p^{2}+x^{2}\right)\right]+1\right)\;.\end{split}$ (194)
@@
$W_{\beta}(x,p)=\dfrac{\left(e^{-([x-\sqrt{2}\beta)^{2}+(p-\sqrt{2}\beta_{i})^{2}]}+e^{-[(x+\sqrt{2}\beta_{r})^{2}+(p+\sqrt{2}\beta_{i})^{2}]}+2e^{-(x^{2}+p^{2})}\cos[2\sqrt{2}(x\beta_{r}-p\beta_{i})-\gamma]\right)}{2\pi\left(1+e^{-2\left|\beta\right|^{2}}\cos\gamma\right)}\;.$ (195)
@@
Appendix F Gaussian quantum Fisher information
@@ F.1 General expression
For mixed Gaussian states in a general Gaussian channel, the QFI reads
$F_{Q}[\hat{\rho}(\theta)]=\frac{1}{2}\frac{\,\mathrm{Tr}\left[(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}])^{2}\right]}{1+\mathcal{P}(\hat{\rho}(\theta))^{2}}+\frac{2\left(\frac{\partial\mathcal{P}(\hat{\rho}(\theta))}{\partial\theta}\right)^{2}}{1-\mathcal{P}(\hat{\rho}(\theta))^{4}}+\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{T}\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right),$ (196)
with the Gaussian purity $\mathcal{P}(\hat{\rho})=(2\sqrt{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]})^{-1}$. This expression can be further simplified using the following properties of $2\times 2$ matrices:
$A^{-1}$ $=\frac{1}{\det{A}}\left[(\,\mathrm{Tr}{A}){I}-{A}\right],$
$(\,\mathrm{Tr}{A}){I}-{A}$ $=\Omega A^{T}\Omega^{T},$
$\,\mathrm{Tr}{\left[A^{2}\right]}$ $=(\,\mathrm{Tr}{A})^{2}-2\det{A},$
$\det A$ $=\frac{1}{2}\,\mathrm{Tr}{\left[A\Omega A^{T}\Omega^{T}\right]},$ (197)
where the last equality reduces to $\mathrm{Tr}{\left[(\Omega H)^{2}\right]}=-2\det H$ for $H^{T}=H$. With these identities, we obtain
$\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}$] $=\frac{\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\Omega^{T}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$, (198)
and
$\,\mathrm{Tr}\left[(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}])^{2}\right]=\,\mathrm{Tr}\left[\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right]^{2}-2\frac{\det\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}}{\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}$. (199)
Moreover, using
[…]
the first and second term in Eq. (196) simplify to
$\frac{1}{2}\frac{\,\mathrm{Tr}\left[(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}])^{2}\right]}{1+\mathcal{P}(\hat{\rho}(\theta))^{2}}=2\frac{\frac{1}{\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}\left(\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right)^{2}-2\det\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}}{1+4\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}$, (201)
and
$\frac{2\left(\frac{\partial\mathcal{P}(\hat{\rho}(\theta))}{\partial\theta}\right)^{2}}{1-\mathcal{P}(\hat{\rho}(\theta))^{4}}=2\frac{\frac{1}{\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}\left(\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right)^{2}}{16\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]^{2}-1},$ (202)
respectively, leading to the Gaussian QFI for single-mode Gaussian channels
$F_{Q}[\hat{\rho}(\theta)]=\frac{8\left(\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right)^{2}}{16\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]^{2}-1}-\frac{4\det\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}}{4\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]+1}+\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{T}\frac{\Omega\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\Omega^{T}}{\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right).$ (203)
For a general Gaussian channel (II.3), the above expression is determined by
$\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}$ $=\frac{\partial\mathcal{X}(\theta)}{\partial\theta}\langle\hat{\bm{r}}\rangle_{\hat{\rho}}+\frac{\partial\bm{d}(\theta)}{\partial\theta}$
$\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}$] $=\mathcal{X}(\theta)\Gamma[\hat{\rho},\mathbf{\hat{r}}]\mathcal{X}(\theta)^{T}+\mathcal{Y}(\theta)$
$\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}$] $=\frac{\partial\mathcal{X}(\theta)}{\partial\theta}\Gamma[\hat{\rho},\mathbf{\hat{r}}]\mathcal{X}(\theta)^{T}+\mathcal{X}(\theta)\Gamma[\hat{\rho},\mathbf{\hat{r}}]\frac{\partial\mathcal{X}(\theta)}{\partial\theta}^{T}+\frac{\partial\mathcal{Y}(\theta)}{\partial\theta}.$ (204)
These terms can be further simplified using Eq. (F.1).
F.2 Applications
F.2.1 Unitary symplectic evolution
A special case of the Gaussian channel is the unitary Gaussian evolution that is described by a symplectic matrix $\mathcal{X}(\theta)=\mathcal{S}(\theta)=\exp(\Omega H\theta)$ and $\mathcal{Y}(\theta)=0$, with no displacement, $\bm{d}(\theta)=0$; see Eq. (II.3). In this case, Eqs. (F.1) read,
$\frac{\partial}{\partial\theta}\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta)}$ $=\Omega H\mathcal{S}(\theta)\langle\hat{\bm{r}}\rangle_{\hat{\rho}}$,
$\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}$] $=\mathcal{S}(\theta)\Gamma[\hat{\rho},\mathbf{\hat{r}}]\mathcal{S}(\theta)^{T},$
@@ F.2.1 Unitary symplectic evolution
In particular, $\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]=0$. Furthermore, we use Eqs. (199) together with $\mathrm{Tr}\left[\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\right]=0$ and the properties (F.1) to show that, in the unitary single-mode case,
@@
$\det\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}=4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\det H-\,\mathrm{Tr}{\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\right]}^{2}.$ (206)
With this, the QFI (203) becomes independent of $\theta$ and reduces to
$F_{Q}[\hat{\rho},\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}$] $=\frac{4\,\mathrm{Tr}\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\right]^{2}-16\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\det{H}}{4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+1}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$. (207)
F.2.2 Symplectic evolution with thermal noise
To account for the effect of diffusive Gaussian noise, caused by weak coupling to a thermal bath, we consider the channel (II.3), described by the operators
$\mathcal{X}(\theta,t)$ $=e^{-\kappa t/2}\mathcal{S}(\theta),$
$\mathcal{Y}(t)$ $=(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\mathbb{I},$
$\bm{d}(\theta)$ $=0,$ (208)
where $\kappa$ describes the coupling to the bath that has an average number of $n_{b}$ excitations. In this model, we can interpret the parameter $\theta=\omega t$ as the product of a frequency factor in front of the Hamiltonian and the evolution time $t$, which is consistent with our previous descriptions of unitary evolutions. Equations (F.2.2) describe the evolution of the system in time $t$, whereas for quantum metrology, we are only interested in the evolution with respect to the parameter $\theta$. To make this distinction explicit, we denote the state at any point in the evolution by $\hat{\rho}(\theta,t)$.
Now, Eqs. (F.1) read
$\frac{\partial}{\partial\theta}\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta,t)}$ $=e^{-\kappa t/2}\Omega H\mathcal{S}(\theta)\langle\hat{\bm{r}}\rangle_{\hat{\rho}}$,
$\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-\kappa t}\underbrace{\mathcal{S}(\theta)\Gamma[\hat{\rho},\mathbf{\hat{r}}]\mathcal{S}(\theta)^{T}}_{\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}]}+(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\mathbb{I},$
$\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-\kappa t}\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}$], (209)
where $\hat{\rho}(\theta,0)$ describes the evolved state in the absence of noise. Using the identity for $2\times 2$ matrices $\det(A+B)=\det A+\det B+\,\mathrm{Tr}A\,\mathrm{Tr}B-\,\mathrm{Tr}\left[AB\right]$, we obtain
$\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-2\kappa t}\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+(1-e^{-\kappa t})^{2}\left(n_{b}+\frac{1}{2}\right)^{2}+e^{-\kappa t}(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\,\mathrm{Tr}{\left[\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}]\right]}$, (210)